1- Suppose a perfectly competitive labour market has a demand curve of LD= 120 – 2w and a supply curve of LS= 8w, where W is the wage rate is dollars and L is the quantity of labour in person-hours
A. What are the equilibrium values of the wage and employment?
B. Suppose the government imposed a minimum wage of $14 per hour, now what are the equilibrium values of the wage and employment?
C. Repeat part (a) assuming now that the market is a monopsony?
D. Repeat part (b) assuming now that the market is a monopsony?
1
Expert's answer
2020-04-14T10:02:14-0400

A)

"120-2w=8w"

"10w=120"


"w=12"

"L=8\times12=96"

B)


"L_S>L_D"

"L=120-2\times14=92"

C)



"MRP=MC_L"

"MC_L=w(1+\frac{1}{E_s})"

"\frac {9}{8}w=120-2w"

"w=38.4"

"L=43.2"

D)




"l_{MIN}=14"

"L=8 \times14=112"